OpenStudy (anonymous):

The slope of the line that contains the points (1, -1) and (2, 3) is 4. What is the y-intercept?

OpenStudy (anonymous):

y-3 =4 --- x-2 y-3=4x-8 y=4x-5 y intercept=-5 u should reward

OpenStudy (anonymous):

how

OpenStudy (anonymous):

money is always nice. other than that, click on "good answer"

OpenStudy (anonymous):

click on good answer